Find the following probabilities based on the standard normal variable Z. (You may find it useful to reference the z table. Leave no cells blank - be certain to enter "0" wherever required. Round your answers to 4 decimal places.)Find the following probabilities based on the standard normal variable Z. (You may find it useful to reference the z table. L
Answer #1

Solution :

Given that ,

a)P( -1.03\leq z \leq -0.62 )

= P(z \leq -0.62) - P(z \leq -1.03)

=0.2676 - 0.1515 = 0.1161

probability = 0.1161

b)

P( 0.04\leq z \leq 2.13 )

= P(z \leq 2.13) - P(z \leq 0.04)

=0.9834 - 0.5160 = 0.4675

probability = 0.4675

c)

P( -1.42 \leq z \leq 0.07)

= P(z \leq 0.07) - P(z \leq -1.42)

=0.5279 - 0.0778 = 0.4501

probability = 0.4501

d)p(z >3)

= 1-p(z< 3)

= 1-0.9987 = 0.0013

probability = 0.0013